OpenStudy (anonymous):

-5x-12<8 x=-4 (-4,inf) is this right

OpenStudy (lgbasallote):

-5x - 12 < 8 -5x < 8 +12 -5x < 20 X > -4 <---you should write it like that NOT x = -4 anyway (-4, \(\infty\)) is right

OpenStudy (anonymous):

ok thanks so much

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

got one more for you to check 4x+3+x>1+2x+8 I got x=2 \[(\infty,2)\]

OpenStudy (anonymous):

igbasallote thanks for the medal

OpenStudy (lgbasallote):

like i said before....STOP using EQUAL SIGNS in inequalities.. and also im pretty sure your ordered pair is wrong.. it's impossible for positive infinity to come BEFORE positive 2

OpenStudy (anonymous):

x>8 \[(-\infty,8)\]

OpenStudy (lgbasallote):

let's check 4x+3+x>1+2x+8 5x + 3 > 9 + 2x 3x > 6 x > 2

OpenStudy (anonymous):

I made a typed wrong \[(-\infty,2)\]

OpenStudy (lgbasallote):

it says x is greater than 2...tell me, is -infinity greater than 2?

OpenStudy (anonymous):

so it is \[(2,\infty)\]

OpenStudy (anonymous):

nope

OpenStudy (lgbasallote):

nope what?

OpenStudy (anonymous):

-inf is not larger than 2

OpenStudy (anonymous):

I type kinda slow

OpenStudy (lgbasallote):

so what would be the right ordered pair?

OpenStudy (anonymous):

(2,inf)

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

welcome